Circular No. 8025

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ji IN NGC 3655 Y. L. Qiu and J. Y. Hu, Beijing Astronomical Observatory (BAO),

report the discovery of an apparent supernova on unfiltered images taken with BAO 0.6-m telescope on Nov. 30.84 (mag ∼ 15.1) and Dec. 1.93 UT (mag ∼ 15.1). SN 2002ji is located at α = 11^{h}22^{m}53^{s}.15, δ = +16^{o}35′10′′.0 (equinox 2000.0), which is 22′′.4 west and 14′′.0 south of the nucleus of NGC 3655. An image taken on Apr. 10 showed nothing at this position (limiting mag ∼ 18.0).

V838 MONOCEROTIS E. Giro, S. Desidera, U. Munari, and C. Pernechele, Astronomical

Observatory, Padua; and G. Baume, Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata and Consejo Nacional de Investigaciones Científicas y Técnicas, report: “Polarimetric imaging of V838 Mon was obtained with the Asiago 1.82-m telescope (+ AFOSC) on Nov. 12.01 UT through a Johnson-V filter, covering the whole extension of the light echo in the north-south direction and 18′′ east-west. The central star displays a 3.05-percent (± 0.35) polarization with position angle 159^{o}.7 ± 3^{o}.9. The light echo shows a distributed polarization with a complex pattern which, south of the central star, is reminiscent of a dipole structure, reaching a 45-percent maximum and 178^{o} position angle at 11′′ south and 3′′ of the central star. The map can be downloaded at

http://www.pd.astro.it/Asiago/IAUC/polmapV838mon.jpg.”

COMETS C/2002 U6, C/2002 U7, C/2002 U8, C/2002 U9 (SOHO) Further to IAUC 8022, D. Hammer reports measurements for comets

found on SOHO website C2 coronagraph images by R. Kracht (C/2002 U6), R. Matson (C/2002 U7), X. Leprette (C/2002 U8), and S. Hönig (C/2002 U9). The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2002-W42.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 U6 Oct. 28.076  14^{h}04.^{m}0  −11^{o}44′
C/2002 U7  30.826  14 13.9  −15 03
C/2002 U8  31.338  14 16.1  −15 16
C/2002 U9  31.435  14 16.0  −15 12
2002 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green